samantharichmond:

If a 60–watt bulb is used to light a room, how much current is flowing through the bulb given that the voltage is 120 volts?

Razor:

Do you still require help for this?

Razor:

P = VI 60 = 120*I I = 60/120 I = 0.5 A

samantharichmond:

no but thank you

Razor:

I hope you stay though, welcome to QC.

samantharichmond:

thank you